OpenStudy (anonymous):

Prove that (1 + x)^n > or equal to (1 + nx) , for all natural number n , where x>-1

OpenStudy (anonymous):

@ajprincess @FoolForMath @siddhantsharan @satellite73 @experimentX

OpenStudy (anonymous):

Mathamatical Induction!

OpenStudy (anonymous):

try induction

OpenStudy (anonymous):

not getting!

OpenStudy (anonymous):

ok did you do the base step?

OpenStudy (anonymous):

yes P(1)

OpenStudy (anonymous):

then assume p(k) is true

OpenStudy (anonymous):

ok so what do we get to assume? that \[(1+x)^k\geq 1+kx\] and then we need to show that given this fact, we can conclude that \[(1+x)^{k+1}\geq 1+(k+1)x\] right?

OpenStudy (anonymous):

we didn't do anything yet, i just wanted to make sure this is the track we are taking

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the reason these are confusing is that it is not an equality, so it is not clear how to manipulate the expression to get what you want. the gimmick here is to multiply both sides of the inequality we get to assume, namely \[(1+x)^k\geq 1+kx\] by \(1+x\) and we get to keep the senses of the inequality because \(x>-1\) so \(1+x>0\)

OpenStudy (anonymous):

how would you know this step? i don't know the answer, but it is the quickest way to get \((1+x)^{k+1}\) out of \((1+x)^k\)

OpenStudy (anonymous):

yes i knw it!!

OpenStudy (anonymous):

would you like to try it yourself, or would you like me to continue?

OpenStudy (anonymous):

i have tried this and i am unable u plzz go on!

OpenStudy (anonymous):

oh ok, so given the induction hypothesis, namely that \[(1+x)^k\geq 1+kx\] we multiply by \(1+x\) and get \[(1+x)^k(1+x)\geq (1+kx)(1+x)\] \[\implies (1+x)^{k+1}\geq 1+x+kx+kx^2\] \[\implies (1+x)^{k+1}\geq 1+(k+1)x+kx^2\] and since \(x^2>0, k>1\) we know \(kx^2>0\) so \[(1+x)^{k+1}\geq 1+(k+1)x\] as required

OpenStudy (anonymous):

thxxx

OpenStudy (anonymous):

yw notice that the induction took care of itself, all that was necessary was some algebra

OpenStudy (anonymous):

yup it needs some logical thinking!!!

OpenStudy (anonymous):

Here is a proof for x >0. Use the Binomial Expansion Theorem \[ (1+x)^n = 1 + n x + {n \choose 2} x^2 + \cdots + x^n \ge 1 + n x \]